The length of a rectangle is 3 feet shorter than twice its width. If the area of the rectangle is 170 ft^2, what is the length if the rectangle?
Answer by jorel1380(3719)

L=2W-3
(2W-3)xW=170
2W2-3W-170=0
(2W+17)(W-10)=0
W=-17 or 10
Throwing out the negative answer, we get Width=10', Length=17'..